OpenStudy (anonymous):

is x(x+3) the same as (x+3)?

OpenStudy (zyi6):

no. because in the first one, it would become 2x + 3, which is not equal to x + 3.

hero (hero):

If x = 1, then yes

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

no as x(x+3) = x^2 + 3x, both are equal only when x = 1

OpenStudy (anonymous):

also when x=-3